Will you please help me write an equation for in slope intercept that passes through given point and is parallel to the graph of the given equation.
(1,-3);y+2=4(x-1)
Thank You

Answer by sudhanshu_kmr(1152)


y + 2 = 4x- 4
=> 4x -y -6 = 0
slope of line = 4
equation of required line ...
(y+3) = 4(x-1)
=> y = 4x -7 .....( slope intercept form)